OpenStudy (anonymous):

Give the equivalent numerator if the denominator is (x-2)(x-1)(x-6). 2x² - 10x + 12 2x² - 14x + 12 2x² - 14x - 10

OpenStudy (anonymous):

\[\frac{ 2 }{ x-2 }\]

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (owlcoffee):

\[\frac{ 2 }{( x-2) }\] Analogically as the last problems, we multiply both numerator and denominator by the expression we require as equal denominator: \[\frac{ 2(x-1)(x-6) }{ (x-2)(x-1)(x-6) }\] All you have to do is apply the distributive property on the numerator of this last fraction.

OpenStudy (anonymous):

How?

OpenStudy (owlcoffee):

\[2(x-1)(x-6)\] \[(2x-2)(x-6)\] \[2x^2-12x+2x+12\] \[2x^2 -10x + 12\] And that's the numerator we were looking for.